# Fully quantum stochastic entropy production

### v1

@@ Abstract
Building on the approach of stochastic thermodynamics, we define entropy production for arbitrary quantum processes, using Bayesian tools to capture the inherent subjective element. First, we show that the classical expression for average entropy production involves only comparisons of statistics at the input or output, and thus can be formally generalized by replacing probability distributions by quantum states. To go beyond the mere average, we construct an entropy production operator, that generalizes the value of entropy to the non-commutative case. The operator is Hermitian; its average is non-negative, so that the Second Law takes its usual form $\left\langle\Sigma\right\rangle\geqslant 0$; it formally satisfies the Jarzynski equality; and recovers the classical expression for classical processes. We also present features whose fully quantum version is not a straightforward translation of the classical one: these may point to the core difference between classical and quantum thermodynamics.
@@ I Introduction
For the second definition (Section IV), we construct a Hermitian entropy production operator using suitable states-over-time to describe both the physical process and its Bayesian reverse. The average entropy production is then the expectation value of this operator over the physical process. Notably, this definition automatically satisfies the Jarzynski equality.
@@ IV.1 Definition
These operators, similarly to $\tilde{Q}_{F}$ and $\tilde{Q}_{R}^{\gamma}$, are valid bipartite states, but in addition their dependence on $\rho$ and $\tau$ is linear. In fact, as shown in Appendix A, $Q_{F}$ and $Q_{R}^{\gamma}$ are both well-defined quantum channels acting on $\rho$ and on $\tau$, respectively. More precisely, $Q_{F}$ is equivalent to the complementary channel of $\mathcal{E}$, up to a transpose. Similarly, $Q_{R}^{\gamma}$ is equivalent to the complementary channel of the reverse map $\mathcal{R}_{\mathcal{E}}^{\gamma}$.
@@ IV.3 Fully quantum fluctuation theorems
By contrast, the generalisation of Crooks’ fluctuation theorem shows a difference between classical and quantum fluctuations in the case of non-commuting processes. Since our operator is Hermitian, we can define the values of entropy production as its eigenvalues:
@@
Next we need to define entropy production in the reverse process. The operator is obviously given by exchanging the roles of $Q_{R}^{\gamma}$ and $Q_{F}$. Contrary to the classical case, the identity $\Sigma[Q_{R}^{\gamma},Q_{F}]=-\Sigma[Q_{F},Q_{R}^{\gamma}]$ does not hold in general. The eigenvalues are indeed the same with opposite signs, as can be proved notcing that $e^{\Sigma[Q_{F},Q_{R}^{\gamma}]}=\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}$ and $e^{-\Sigma[Q_{R}^{\gamma},Q_{F}]}=(Q_{R}^{\gamma})^{-1/2}Q_{F}(Q_{R}^{\gamma})^{-1/2}$ have the same eigenvalues by polar decomposition. The eigenvectors are in general different, and so
@@
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,\langle r_{k}|Q_{R}^{\gamma}|r_{k}\rangle=\,\sum_{j}e^{-\Sigma_{j}}|\langle r_{k}|\sqrt{Q_{F}}|f_{j}\rangle|^{2}\,,$ (19)
where we have rewritten Eq. 13 as $Q_{R}^{\gamma}=\sqrt{Q_{F}}e^{-\Sigma[Q_{F},Q_{R}^{\gamma}]}\sqrt{Q_{F}}$. Eq. 19 is the fully quantum extension of the Crooks theorem. The classical Crooks expression
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,e^{-\Sigma_{k}}\,\mathbb{P}_{F}(\Sigma_{k})$ (20)
is recovered only when $[Q_{F},Q_{R}^{\gamma}]=0$. In the general case, non-commutativity (i.e. coherence) impacts the fluctuations of entropy. Still, like in the classical case, Jarzynski equality (IV.3) can be derived starting from Eq. 19, as $\sum_{k}\mathbb{P}_{R}(-\Sigma_{k})=\sum_{j}e^{-\Sigma_{j}}\mathbb{P}_{F}(\Sigma_{j})=1$. A comparison between the quantum and classical expressions is shown in Fig. 1 for the channel that we shall use in Section V. One might surmise that quantum fluctuations are always larger than classical ones. The figure shows that this is often the case, but not always: for some $\rho$, the standard deviation of the quantum distribution is actually smaller.
Figure 1: Comparison between the statistics of entropy in the quantum Eq. 19 and the classical Crooks expression Eq. 20 for a one-qubit channel (the entropy production operator is of dimension 4). The channel $\mathcal{N}$ is defined by Eq. 35 where $\gamma$ is chosen to be equal to $\xi$, and $\tau=\mathcal{N}(\rho)$. For the numerics, we choose $\langle 0|\xi|0\rangle=0.8$, $\phi=0.8$ and $n=1$. The plot represents the $(x,z)$ plane of the Bloch sphere of $\rho$. The heat map represents the standard deviation of the reverse distribution $\mathbb{P}_{R}$: left, for Eq. 19 obtained from studying our operator; right, computed from $\mathbb{P}_{F}$ assuming the classical Crooks relation Eq. 20. For most $\rho$, the quantum expression shows a larger standard deviation than the classical guess; but for some, the opposite holds.
@@ IV.5 Explicit expressions for the average entropy production
The proof is in Appendix B. The equality does not hold in general, but it does hold if we further have $[\mathcal{E}(\gamma),\tau]\equiv[\mathcal{E}(\gamma),\mathcal{E}(\rho)]=0$, because then it follows directly from (24).
@@ IV.6 Superadditivity and mutual entropy production
for any choice of $\tau_{1}$ and $\tau_{2}$. Equality holds if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$ and $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=0$. The proof of this property is in Appendix C.
@@ VI Conclusion
The bulk of the paper was devoted to an approach that was not attempted previously: the construction of an entropy production operator based on quantum versions of the input-output processes (Definition 2). The operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$ that we constructed [Eq. 13] is Hermitian, has non-negative average value, and formally fulfills Jarzynski equality for every process. Its average recovers Definition 1 [Eq. 24] when $[\mathcal{E}(\gamma),\tau]=0$ and the channel is full rank, up to the contribution at the input being measured by the Belavkin-Staszewski relative entropy instead of the more common Umegaki one.
@@
First of all, we were not able to find an entropy production operator, whose average is always given by the very pleasant Definition 1. Perhaps it can be found—or perhaps, the fully quantum average entropy production must differ in general. As a hint that the latter may be the case, we bring up another work of ours, the derivation of quantum Bayesian retrodiction from a minimum change principle [37]. The map derived there coincides with the Petz map when $[\mathcal{E}(\gamma),\tau]=0$: intriguingly, this is the same condition under which average entropy production falls back onto the familiar form. We have considered the possibility of defining the reverse quantum process not with the Petz map, but with the general expression found in that work. This would amount at replacing Eq. 12 with
@@ Appendix A Physical interpretation of $Q_{F}$ and $Q_{R}^{\gamma}$
Appendix B Proof of Eq. 26
@@
Appendix C Superadditivity of average entropy production

### v2

@@ Abstract
Building on the approach of stochastic thermodynamics, we define entropy production for arbitrary quantum processes, using Bayesian tools to capture the inherent subjective element. First, we show that the classical expression for average entropy production involves only comparisons of statistics at the input or output, and thus can be formally generalized by replacing probability distributions by quantum states. To go beyond the mere average, we construct an entropy production operator, that generalizes the value of entropy to the non-commutative case. The operator is Hermitian; its average is non-negative, so that the Second Law takes its usual form $\left\langle\Sigma\right\rangle\geqslant 0$; it formally satisfies the Jarzynski equality and the Crooks fluctuation theorem; and recovers the classical expression for classical processes. We also present features whose fully quantum version is not a straightforward translation of the classical one: these may point to the core difference between classical and quantum thermodynamics.
@@ I Introduction
For the second definition (Section IV), we construct a Hermitian entropy production operator using suitable states-over-time to describe both the physical process and its Bayesian reverse. The average entropy production is then the expectation value of this operator over the physical process. Notably, this definition satisfies both the Jarzynski equality and Crooks’ fluctuation relation.
@@ IV.1 Definition
These operators, similarly to $\tilde{Q}_{F}$ and $\tilde{Q}_{R}^{\gamma}$, are valid bipartite states, but in addition their dependence on $\rho$ and $\tau$ is linear. In fact, as shown in Appendix A, $Q_{F}$ (up to transposition) and $Q_{R}^{\gamma}$ are both well-defined quantum channels acting on $\rho$ and on $\tau$, respectively. More precisely, $Q_{F}$ is equivalent to the complementary channel of $\mathcal{E}$, up to a transpose. Similarly, $Q_{R}^{\gamma}$ is equivalent to the complementary channel of the reverse map $\mathcal{R}_{\mathcal{E}}^{\gamma}$.
@@ IV.3 Fully quantum fluctuation theorems
Next we turn to Crooks fluctuation theorem, a relation between the distributions of values of entropy in the forward and the reverse process. Since our operator is Hermitian, we can define the values of entropy production as its eigenvalues:
@@
Next we need to define the entropy production in the reverse process. The operator is obviously given by exchanging the roles of $Q_{R}^{\gamma}$ and $Q_{F}$. In contrast to the classical case, the operator identity $\Sigma[Q_{R}^{\gamma},Q_{F}]=-\Sigma[Q_{F},Q_{R}^{\gamma}]$ does not hold in general. The eigenvalues are the same with opposite signs, but the eigenvectors are in general different (see Appendix B):
@@
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,\langle r_{k}|Q_{R}^{\gamma}|r_{k}\rangle\,.$ (19)
Although it is not immediately obvious, by carefully studying the singular value decomposition of $\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}$, one can prove (see Appendix B) that the probability distributions just defined in Eqs. 17 and 19 obey the Crooks relation
$\mathbb{P}_{R}(-\Sigma_{k})$ $=\,e^{-\Sigma_{k}}\,\mathbb{P}_{F}(\Sigma_{k})\,.$ (20)
@@ IV.5 Explicit expressions for the average entropy production
The proof is in Appendix C. The equality does not hold in general, but it does hold if we further have $[\mathcal{E}(\gamma),\tau]\equiv[\mathcal{E}(\gamma),\mathcal{E}(\rho)]=0$, because then it follows directly from (24).
@@ IV.6 Superadditivity and mutual entropy production
for any choice of $\tau_{1}$ and $\tau_{2}$. Equality holds if and only if $\tau_{1}=\mathcal{E}_{1}(\rho)$ and $[\mathcal{E}_{1}(\rho),\mathcal{E}_{1}(\gamma)]=0$. The proof of this property is in Appendix D.
@@ VI Conclusion
The bulk of the paper was devoted to an approach that was not attempted previously: the construction of an entropy production operator based on quantum versions of the input-output processes (Definition 2). The operator $\Sigma[Q_{F},Q_{R}^{\gamma}]$ that we constructed [Eq. 13] is Hermitian, has non-negative average value, and formally fulfills Jarzynski equality and the Crooks fluctuation theorem for every process. Its average recovers Definition 1 [Eq. 24] when $[\mathcal{E}(\gamma),\tau]=0$ and the channel is full rank, up to the contribution at the input being measured by the Belavkin-Staszewski relative entropy instead of the more common Umegaki one.
@@
First of all, we were not able to find an entropy production operator, whose average is always given by the very pleasant Definition 1. Perhaps it can be found—or perhaps, the fully quantum average entropy production must differ in general. As a hint that the latter may be the case, we bring up another work of ours, the derivation of quantum Bayesian retrodiction from a minimum change principle [37]. The map derived there coincides with the Petz map when $[\mathcal{E}(\gamma),\tau]=0$: intriguingly, this is the same condition under which the average entropy production described here assumes the familiar form. We have considered the possibility of defining the reverse quantum process not with the Petz map, but with the general expression found in that work. This would amount at replacing Eq. 12 with
@@ Appendix A Physical interpretation of $Q_{F}$ and $Q_{R}^{\gamma}$
Appendix B Proof of quantum Crooks fluctuation theorem
Here we prove Eq. 20 with the entropy production defined in Eqs. 17 and 19.
Assume $Q_{F}$ and $Q_{R}^{\gamma}$ to be full rank. Let the singular value decomposition of $\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}$ be
$\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1/2}=USV^{\dagger}\,,$ (69)
where $U$ and $V$ are unitary operators, and $S$ is a positive definite diagonal operator. They can be written as
$U=\sum_{k}|f_{k}\rangle\!\!\!\;\langle k|,\quad V=\sum_{k}|r_{k}\rangle\!\!\!\;\langle k|,\quad S=\sum_{k}s_{k}|k\rangle\!\!\!\;\langle k|,\quad s_{k}>0$ (70)
for some orthonormal basis {|k}$\{\ket{k}\}$. {|$\{\ket{f_{k}}\}$ and {|$\{\ket{r_{k}}\}$ are the eigenvectors of $\Sigma[Q_{F},Q_{R}^{\gamma}]$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]$, which can be seen from
$e^{\Sigma[Q_{F},Q_{R}^{\gamma}]}$ $=\sqrt{Q_{F}}(Q_{R}^{\gamma})^{-1}\sqrt{Q_{F}}=US^{2}U^{\dagger}$ (71)
$e^{-\Sigma[Q_{R}^{\gamma},Q_{F}]}$ $=(Q_{R}^{\gamma})^{-1/2}Q_{F}(Q_{R}^{\gamma})^{-1/2}=VS^{2}V^{\dagger}$ (72)
and therefore, $\Sigma[Q_{F},Q_{R}^{\gamma}]=2U(\log S)U^{\dagger}$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]=-2V(\log S)V^{\dagger}$. This also verifies that $\Sigma[Q_{F},Q_{R}^{\gamma}]$ and $\Sigma[Q_{R}^{\gamma},Q_{F}]$ have opposite eigenvalues and $\Sigma_{k}=2\log s_{k}$.
From Eq. 69 and that $Q_{F},Q_{R}^{\gamma}$ are Hermitian, one gets $Q_{F}=USV^{\dagger}Q_{R}^{\gamma}VSU^{\dagger}$. From Eq. 70, one gets |$\ket{f_{k}}=UV^{\dagger}\ket{r_{k}}$. From these, for any $k$,
$\mathbb{P}_{F}(\Sigma_{k})$ $=\bra{f_{k}}Q_{F}\ket{f_{k}}$ (73)
$=\bra{r_{k}}VU^{\dagger}\times USV^{\dagger}Q_{R}^{\gamma}VSU^{\dagger}\times UV^{\dagger}\ket{r_{k}}$ (74)
$=\bra{r_{k}}VSV^{\dagger}Q_{R}^{\gamma}VSV^{\dagger}\ket{r_{k}}$ (75)
$=s_{k}^{2}\bra{r_{k}}Q_{R}^{\gamma}\ket{r_{k}}$ (76)
$=s_{k}^{2}~{}\mathbb{P}_{R}(-\Sigma_{k})$ (77)
noting that |$\ket{r_{k}}$ is the eigenvector of $VSV^{\dagger}$ and $VSV^{\dagger}\ket{r_{k}}=s_{k}\ket{r_{k}}$. Last, since $\Sigma_{k}=2\log s_{k}$, one gets
$\mathbb{P}_{R}(-\Sigma_{k})=e^{-\Sigma_{k}}\mathbb{P}_{F}(\Sigma_{k})$ (78)
which is the Crooks fluctuation theorem in Eq. 20.
Appendix C Proof of Eq. 26
@@ Appendix B Proof of quantum Crooks fluctuation theorem
Appendix D Superadditivity of average entropy production